A mixture of ethane and ethene occupies 41L at 1 atm and 500 K. The mixture reacys completely with 10/3 mole of O2 to produce CO2 and H2O. The mole fractions of ethane and ethene in the mixture are (R= 0.082 L atm/K mol)a) 0.50,0.50b) 0.75,0.25c) 0.67,0.33d) 0.25,0.25e) 0.33,0.67

Grade:12

1 Answers

Sanketh
6 years ago
PV=nRT
1×41=n0.082×500
n=1 =》n1+n2=1  {1}
C2h6+7/2o2-》2co2+3h2o
C2h4+3o2-》2co2+2h2o
7/2n1+3n2=10/3  {2}
Solving {1} n {2} 
n1=2/3 and n2=1/3
X1=2/3
=0.67
X2=1/3
=0.33
Ans:c